OpenStudy (anonymous):

Factor the Following Polynomials: x²-81 x²-13x+36 Which polynomial is a special product and why?

OpenStudy (anonymous):

could you walk me through the work process

OpenStudy (anonymous):

(x + y)(x − y) = x^2 − y^2 is a special product x^2-81=(x+9)(x-9),this is a special product x^2-13x+36 x^2-4x-9x+36 (x-4)(x-9)